integrate the following integral
∫ cos^2(nx) dx 

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

I = \int cos^{2}(nx)dx
I =\frac{1}{2} \int 2cos^{2}(nx)dx
I =\frac{1}{2} \int (1+cos(2nx)dx
I =\frac{1}{2}[\int dx+\int cos(2nx)dx]
I =\frac{1}{2}[x+\int cos(2nx)dx]
I =\frac{1}{2}[x+\frac{sin(2nx)}{2n}]+c